Find the average value gave of the function g on the given interval. g(x) = 77, [27, 64] Save =
Answer #1

Solution

Average value of any function \bar{f} =\frac{1}{b-a}\int_{a}^{b}f(x)dx

g(x)=7\sqrt[3]{x} [27, 64]

\bar{g}(x) =\frac{1}{64-27}\int_{27}^{64}g(x)dx

\bar{g}(x) =\frac{1}{37}\int_{27}^{64}7\sqrt[3]{x}(dx)

\bar{g}(x) =\frac{7}{37}\int_{27}^{64}x^{\frac{1}{3}}(dx)

\bar{g}(x) =\frac{7}{37}\times \frac{3}{4}[x^{\frac{4}{3}}]^{64}_{27}

\bar{g}(x) =\frac{21}{148}[256-81]

\bar{g}(x) =\frac{21}{148}\times [175]

\bar{g}(x) =\frac{3675}{148}

\bar{g}(x) =24.83
Answer #2

SOLUTION :


Average value ifbtwen given function g(x) in interval (27, 64), gavg


   64

= ∫    g(x) dx  / (64 - 27)

   27


              64

= 1/37   ∫    7 (x^(1/3)  dx

              27

                                 64

= 7/37 * 3/4 [ t^4/3) ]                            

                                27                                                       


= 21/148 [64^(4/3) - 27^(4/3)]


= 21/148 * 175


= 24.83 (ANSWER)

answered by: Tulsiram Garg